a parallel combination of three resistors takes a current of 7.5A from 30 V supply.if the two resistors are 10 ohm and 12 ohm, find the third one.
1
Expert's answer
2012-07-19T07:37:13-0400
The resistance of the group of 3 resistors is equal to

R = 30 V / 7.5 A =
4 ohm.

Let R3 be the resistance of the third resistor, then

1/R =
1/10 + 1/12 + 1/R3

so,

1/4 = 1/10 + 1/12 + 1/R3

1/R3 = 1/4
- 1/10 - 1/12
= 15/60 - 6/60 - 5/60
= 4/60

whence


R3 = 60/4 = 15 ohm